OpenStudy (anonymous):

Geometric sum. Should be a fun one :). Evaluate the sum 1/2 + 1/6 + 1/18 + 1/54 + ... + 1/1458

OpenStudy (anonymous):

I know how to put other problems like this into an equation to get the sum. The difference with this one is it's not telling me how many terms there are in between.

OpenStudy (anonymous):

I do realize somewhere along the lines it's multiplying each term by 1/3

OpenStudy (anonymous):

Geometric sum equation \[a1(\frac{1-r^n }{1-r})\]

OpenStudy (anonymous):

I agree with that :)

OpenStudy (anonymous):

Problem is need to find n I think.

OpenStudy (anonymous):

a1=1st term r=your common ration (1/3 like you said) n=number of terms

OpenStudy (anonymous):

it contain total 7 terams since (3^6)*2= 1458 so find sum using the eqn

OpenStudy (anonymous):

Well thats where we go to the geometric sequence formula: \[a _{n}=a_{1}r^{n-1}\]

OpenStudy (anonymous):

\[a _{n}\]=the last term in the sequence and we already know that to be 1458 and all the other variables are the same like in the last formula

OpenStudy (anonymous):

the answer is 1093/1458

OpenStudy (anonymous):

So I see how you plugged it in. We need to solve for that n right?

OpenStudy (anonymous):

So you end up with log2(1/729) = n-1 ?

OpenStudy (anonymous):

No, sorry my bad

OpenStudy (anonymous):

Hold up I put the worng number up \[\frac{ 1 }{ 1458 }=\frac{ 1 }{ 2 }*\frac{ 1 }{ 3 }^{(n-1)}\]

OpenStudy (anonymous):

It's ok :), take your time. I appreciate the help.

OpenStudy (anonymous):

\[\frac{ 1 }{ 729 }=\frac{ 1 }{ 3 }^{n-1}\]

OpenStudy (anonymous):

This makes makes more sense to me now, lol

OpenStudy (anonymous):

\[\log_{\frac{1}{3}} \frac{ 1 }{ 729 }=\log_{\frac{1}{3}}\frac{1}{3}^{n-1}\]

OpenStudy (anonymous):

\[6=n-1\] \[7=n\]

OpenStudy (anonymous):

OK we got our "n" now we can proceed :)

OpenStudy (anonymous):

very cool :).

OpenStudy (anonymous):

@Jhannybean Its called the change of base formula

OpenStudy (anonymous):

Ok, lol that ws easy anywyas back to @bigmatt \[\frac{1}{2}(\frac{1-1/3^{7}}{1-(1/3)})\]

OpenStudy (anonymous):

We get 1093/1458 I think.

OpenStudy (anonymous):

That's it, lol I good, I didn;t want to compute that, unless you don't see it, anything else?

OpenStudy (anonymous):

Ya, that's it! I just needed help with finding n and you did great :) Thank you!

OpenStudy (anonymous):

Np, anytime